OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form. answer choices: –2x – y = –7 2x – y = –7 2x – y = 11 –2x + y = –11

OpenStudy (helder_edwin):

u have a point (x0,y0) and the slope m. so u can use \[ \large y-y_0=m(x-x_0) \]

OpenStudy (anonymous):

I just don't know how to write an equation for it... @helder_edwin

OpenStudy (helder_edwin):

ok u have m=2 and have the point (x0,y0)=(-1,5) then \[ \large y-5=2(x-(-1)) \]

OpenStudy (helder_edwin):

this becomes \[ \large y-5=2(x+1) \] \[ \large y-5=2x+2 \] \[ \large -5-2=2x-y \] \[ \large 2x-y=-7 \]

OpenStudy (helder_edwin):

got it?

OpenStudy (anonymous):

Yes @helder_edwin